Two light bulbs, have exponential lifetime where expected lifetime for bulb A is 500 hours and expected lifetime for bulb B is 200 hours.

a) What is the expected time until bulb A or bulb B malfunctions ?

b) What is the probability that bulb A malfunctions before bulb B ?  
Answer #1

a)here parameter \lambdaA =1/500 and \lambdaB =1/200

a) expected time until bulb A or bulb B malfunctions=1/(\lambdaA +\lambdaB ) =1/(1/500+1/200)=142.8571 Hours

b)

probability that bulb A malfunctions before bulb B =P(A<B)=\lambdaA/(\lambdaA+\lambdaB)=(1/500)/(1/500+1/200)

=0.2857